place the following fractions 3/5, 1/8 and 2/7 on a number line.

Answer by Theo(13342)

1/8 is less than 2/7.
2/7 is less than 3/5.

the fractions should be in the following order from left to right on the number line.

1/8, 2/7, 3/5.

1/8 * 7/7 = 7/56
2/7 * 8/8 = 16/56
16/56 is greater than 7/56.

2/7 * 5/5 = 10/35
3/5 * 7/7 = 21/35
21/35 is greater than 10/35.

the order on the number line is correct.

1/8 is to the left of 2/7 which is to the left of 3/5.